# Abelian surfaces over finite fields containing no curves of genus $3$ or less

### v1

@@ 3. Factorisation of $2$
Let $f(t)$ be a polynomial from $\mathcal{P}^{\mathrm{irr}}_{\mathrm{npp}}\sqcup\mathcal{P}^{\mathrm{irr}}_{\mathrm{Wres}}$. For the rest of the paper, we set $K=\mathbb{Q}[t]/f(t)$. Then $K$ is a CM-field, that is, a totally imaginary quadratic extension of a totally real field $K^{+}$. Concretely, if we denote by $\pi$ the class of $t$ in $K$, then $K^{+}=\mathbb{Q}(\pi+q/\pi)$. Moreover, if we write
@@
Since $f(t)\in\mathcal{P}^{\mathrm{irr}}_{\mathrm{Wres}}$ then $a=0$. Hence, $\Delta_{f^{+}}=4(2q-b)$. Let $d$ be a positive squarefree integer such that $2q-b=c^{2}d$ with $c\in\mathbb{Z}$. Note that if $c$ is odd then $2q-b\equiv d\bmod 4$. Recall that $2$ ramifies in $K^{+}=\mathbb{Q}(\sqrt{d})$ if and only if $d=1$ or $d\equiv 2,3\bmod 4$. We now analyse all the possible cases, which are described in Theorem 1.2.(3b).
@@
Clearly, ii $\Rightarrow$ i and iii $\Rightarrow$ i. Before proving the remaining implications, we pause to prove a general claim. Set $N:=N_{K^{+}/\mathbb{Q}}(\Delta_{K/K^{+}})$. We have
@@
Assume first that $f$ is non-ordinary. Then $f(t)=t^{4}-qt^{2}+q^{2}$ by Theorem 1.2 and Lemma 1.9. Note that $\zeta_{3}\mapsto-\pi^{2}/q$ gives an embedding of $\mathbb{Q}(\zeta_{3})$ in $K$ (cf. [14, Sec. 4]). Moreover, as we have shown in the proof of Lemma 3.1, we have that $K^{+}=\mathbb{Q}(\sqrt{3})$ or $K^{+}=\mathbb{Q}(\sqrt{6})$. In the first case, we have $K=\mathbb{Q}(\zeta_{3},\sqrt{3})$ and one computes that $\Delta_{K^{+}}=12$ and $\Delta_{K}=12^{2}$. In the second case, we have $K=\mathbb{Q}(\zeta_{3},\sqrt{6})$ and one computes that $\Delta_{K^{+}}=24$ and $\Delta_{K}=24^{2}$. In both cases $K/K^{+}$ is unramified.
Hence, $f(t)$ must be ordinary. Then, by Lemma 1.9, either $b=1-2q$, or $b=2-2q$ and $q$ is odd. In the first case, Equation (1) implies that $N=1$, which cannot happen by assumption. So we must be in the second case, that is, iv holds. Equation (1) implies that if $b=2-2q$ (with $q$ odd) then $N\in\{1,2,4\}$, and in fact $N\in\{2,4\}$ since we are assuming that $K/K^{+}$ is ramified. Since $\mathfrak{m}$ is the unique maximal ideal of $\mathcal{O}^{+}$ above $2$ by Lemma 3.1, we obtain that $\mathfrak{m}$ ramifies in $K$, that is, ii holds.
To conclude, we prove that iv $\Rightarrow$ iii. Since $b=2-2q$ with $q$ odd, we have that $f(t)\equiv(t+1)^{4}\bmod 2$. By the Kummer-Dedekind theorem [24, Thm. 8.2], the order $\mathbb{Z}[\pi]$ in $K$ has a unique maximal ideal above $2$ which is regular since the remainder of the division of $f(t)$ by $t+1$ is $(q-1)^{2}+2\equiv 2\bmod 2^{2}$, because $q$ is odd. This implies that $\mathcal{O}$ has a unique maximal ideal above $2$ with residue field $\mathbb{F}_{2}$, that is, $2$ is totally ramified in $K$. ∎
@@
If $K/K^{+}$ is ramified or if there exists a prime in $K^{+}$ that divides $(\pi-q/\pi)$ and which stays inert in $K$, then, by [13, Thm. 1.1], there is a principally polarised abelian surface in the isogeny class associated to $f(t)$. Hence $f\in\mathcal{P}^{\mathrm{irr}}_{\mathrm{Wres}}$ by Theorem 1.2.
@@
If $f(t)$ is ordinary then [12, proof of Lemma 12.2] shows that $K$ contains $M=\mathbb{Q}(\zeta_{3})$. If $f(t)$ is not ordinary, then [14, Thm. 2] tells us that
@@
Then $\zeta_{3}\mapsto-\pi^{2}/q$ gives an embedding of $M$ in $K$ (cf. [14, Sec. 4]). ∎
@@
Let $f\in\mathcal{P}^{\mathrm{irr}}_{\mathrm{npp}}\sqcup\mathcal{P}^{\mathrm{irr}}_{\mathrm{Wres}}$. The possible factorisations of $2$ in $K$ are as follows and all of them do occur.
@@
By Lemma 1.11, $2$ is not inert in $K$. We now show that $2$ cannot be totally split in $K$. If $f\in\mathcal{P}^{\mathrm{irr}}_{\mathrm{Wres}}$ then $2$ ramifies in $K^{+}$ by Lemma 3.1 and we are done. Assume then that $f\in\mathcal{P}^{\mathrm{irr}}_{\mathrm{npp}}$. Then, by Lemma 3.4, $K$ contains $M=\mathbb{Q}(\zeta_{3})$. Note that $2$ is inert in $M$, so $2$ is not totally split in $K$.
@@
We start by proving i. So, assume that $f\in\mathcal{P}^{\mathrm{irr}}_{\mathrm{npp}}$. Then $M=\mathbb{Q}(\zeta_{3})$ is a subfield of $K$ by Lemma 3.4. Since $2$ is inert in $M$ then $2$ cannot be totally ramified in $K$. Also, $K/K^{+}$ is unramified by Lemma 3.3. If $2$ is inert in $K^{+}$ then we must have $2\mathcal{O}=\mathfrak{M}_{1}\mathfrak{M}_{2}$, with $\mathfrak{M}_{1}=\bar{\mathfrak{M}}_{2}$. If $2$ splits in $K^{+}$ then $2\mathcal{O}=\mathfrak{m}_{1}\mathfrak{m}_{2}$, that is each maximal ideal of $O^{+}$ above $2$ stays inert in $K$. In particular, $\mathfrak{M}_{1}=\overline{\mathfrak{M}}_{1}$ and $\mathfrak{M}_{2}=\overline{\mathfrak{M}}_{2}$. If $2$ is ramified in $K^{+}$ then $2\mathcal{O}=\mathfrak{M}^{2}$ must occur and $\mathfrak{M}$ is stable under conjugation. Part ii is Lemma 3.1. A search on the Weil polynomials shows that all listed factorisation occur; see Example 6.4. ∎
@@ 4. Kernels of polarisations of abelian varieties
$\mathcal{B}(S):=\frac{Z(S)}{(B(S)\cdot\mathop{\rm Pr}\nolimits_{S}(K^{\dagger}))},$
@@
    $f^{+}\bmod\ell$ has a linear factor.
@@
We now show that i implies iii. So, assume also that $M$ as above satisfies $|M|=\ell^{2}$ for a rational prime $\ell$. Then $r=1$, $n_{1}=1$ and $R^{+}/\mathfrak{n}_{1}\simeq\mathbb{F}_{\ell}$. This implies that $f^{+}\bmod\ell$ has a linear factor (cf. [24, Thm. 8.2]).
@@
This is part of [13, Prop. 6.2]. ∎
@@
This is a special case of [13, Prop. 6.4]. ∎
@@
This is part of [13, Prop. 7.1]. ∎
@@ 5. Polarisations of degree $4$ on abelian surfaces
In Section 2, we showed that an abelian surface with Weil polynomial $f\in\mathcal{P}^{\mathrm{irr}}_{\mathrm{npp}}\sqcup\mathcal{P}^{\mathrm{irr}}_{\mathrm{Wres}}$ contains curve of arithmetic genus $3$ if and only if it admits a polarisation with kernel of order $4$. We now apply the results developed in Section 4 to characterise isogeny classes defined by a polynomial in $\mathcal{P}^{\mathrm{irr}}_{\mathrm{npp}}\sqcup\mathcal{P}^{\mathrm{irr}}_{\mathrm{Wres}}$ not containing abelian surfaces with a curve of genus $3$ lying on.
@@
Since $f\in\mathcal{P}^{\mathrm{irr}}_{\mathrm{Wres}}$, the obstruction element $I_{\mathcal{C}}$ is trivial by Theorem 4.10. Hence, there is an $R$-module $M$ of order $4$ which is attainable if and only if the image $z_{\mathfrak{l}}$ of $[R/\mathfrak{l}R]_{R}$ in $\mathcal{B}(R)$ is trivial. Note that
@@
In Theorem 5.7, we wrap-up the discussion above. Moreover, in certain cases with $\mathcal{C}$ ordinary, we show that if there is an abelian variety with polarisation with kernel of order $4$ then there is an abelian variety with polarisation with kernel of order $4$ and maximal endomorphism ring. To do so, we use results from [12]. The construction of $\mathcal{B}(\mathcal{O})$ in [12] is slightly different form the one in [13] that we have presented above, but certainly equivalent in view of Proposition 4.8 and [12, Prop. 10.1].
@@
Clearly strongly attainable implies attainable, while the contrary does not hold in general. A characterisation of strongly attainable $R$-modules is given in [12, Prop. 5.7], repeated below as Proposition 5.5 for convenience. Such characterisation relates modules with a particular element $I_{K,\Phi}$ of $\mathcal{B}(\mathcal{O})$ which depends on a choice of a CM-type $\Phi$ of $K$. The definition of $I_{K,\Phi}$ is given in [12, Def. 5.2]. For our purposes, it sufficient to know whether it is trivial or not in $\mathcal{B}(\mathcal{O})$.
@@
Assume that $f\in\mathcal{P}^{\mathrm{irr}}_{\mathrm{npp}}$ is ordinary. Then $\mathcal{C}$ does not contain a principally polarised abelian variety. The result follows from [12, Prop. 11.3] and [12, Cor. 11.4].
Assume from now that $f\in\mathcal{P}^{\mathrm{irr}}_{\mathrm{Wres}}$ is ordinary. If $K/K^{+}$ is ramified then the $I_{K,\Phi}=0$ by [12, Prop. 11.1]. If $K/K^{+}$ is unramified and there is no maximal ideal of $\mathcal{O}^{+}$ dividing $\pi-q/\pi$ which stays inert in $K/K^{+}$ then the result follows, again, by combining [12, Prop. 11.3] and [12, Cor. 11.4] and the observation that $\mathcal{C}$ contains a principally polarised abelian variety. ∎
@@
[12, Prop. 5.7] Let $M$ be finite $R$-module. Then $M$ is strongly attainable if and only if $M\simeq\mathcal{O}/\mathfrak{a}\mathcal{O}$ for some $\mathcal{O}^{+}$-ideal $\mathfrak{a}\subseteq\mathcal{O}^{+}$ such that the image in $\mathcal{B}(\mathcal{O})$ of the class of $\mathcal{O}^{+}/\mathfrak{a}$ in $G(\mathop{\rm Mod}\nolimits_{\mathcal{O}^{+}})$ equals $I_{K,\Phi}$.
@@
Assume that $f\in\mathcal{P}^{\mathrm{irr}}_{\mathrm{npp}}\sqcup\mathcal{P}^{\mathrm{irr}}_{\mathrm{Wres}}$.
@@
In the case when $f(t)\in\mathcal{P}^{\mathrm{irr}}_{\mathrm{Wres}}$, we completely characterise when the equivalent conditions ia,ib and iiif hold in terms of the coefficients of $f(t)$ in iiig. It is easy to obtain a characterisation in terms of the coefficients of $f(t)=t^{4}+at^{3}+bt^{2}+aqt+q^{2}$ also when $f(t)\in\mathcal{P}^{\mathrm{irr}}_{\mathrm{npp}}$. Indeed, if we write $\Delta_{f^{+}}=a^{2}-4(b-2q)=c^{2}d$ for integers $c$ and $d$ with $d$ squarefree, then it is well known that iie holds if and only if $d\equiv 5\bmod 8$. Moreover, if $q$ is even (or equivalently $a$ is odd) then $f^{+}(t)\equiv t^{2}+t+1\bmod 2$ is irreducible, which implies that $2$ is inert in $K^{+}$. Similarly, if $a$ is even and $a+b\not\equiv 1\bmod 4$ then $f^{+}(t)\equiv(t+1)^{2}$ and the remainder of the division of $f^{+}(t)$ by $t+1$ is not divisible by $4$. Hence, $2$ ramifies in $K^{+}$ by the Kummer–Dedekind Theorem [24, Thm. 8.2].

### v3

@@ 3. Factorisation of $2$
To prove Main Theorem 3, we first need some technical results concerning the factorisation of $2$ in the extension $K=\mathbb{Q}[t]/f(t)$ where $f(t)$ is a Weil polynomial in $\mathcal{P}^{\mathrm{irr}}_{\mathrm{npp}}\sqcup\mathcal{P}^{\mathrm{irr}}_{\mathrm{Wres}}$. Note that $K$ is a CM-field, that is, a totally imaginary quadratic extension of a totally real field $K^{+}$. Concretely, if we denote by $\pi$ the class of $t$ in $K$, then $K^{+}=\mathbb{Q}(\pi+q/\pi)$. Moreover, if we write
@@
Since $f(t)\in\mathcal{P}^{\mathrm{irr}}_{\mathrm{Wres}}$ then $a=0$. Hence, $\Delta_{f^{+}}=4(2q-b)$. Let $d$ be a positive squarefree integer such that $2q-b=c^{2}d$ with $c\in\mathbb{Z}$. Note that if $c$ is odd then $2q-b\equiv d\bmod 4$. Recall that $2$ ramifies in $K^{+}=\mathbb{Q}(\sqrt{d})$ if and only if $d=1$ or $d\equiv 2,3\bmod 4$. We now analyse all the possible cases, which are described in Theorem 1.4.(iii)(iii).(b).
@@
[…]
@@
Assume first that $f$ is non-ordinary. Then $f(t)=t^{4}-qt^{2}+q^{2}$ by Theorem 1.4 and Lemma 1.12. Note that $\zeta_{3}\mapsto-\pi^{2}/q$ gives an embedding of $\mathbb{Q}(\zeta_{3})$ in $K$ (cf. [15, Sec. 4]). Moreover, as we have shown in the proof of Lemma 3.1, we have that $K^{+}=\mathbb{Q}(\sqrt{3})$ or $K^{+}=\mathbb{Q}(\sqrt{6})$. In the first case, we have $K=\mathbb{Q}(\zeta_{3},\sqrt{3})$ and one computes that $\Delta_{K^{+}}=12$ and $\Delta_{K}=12^{2}$. In the second case, we have $K=\mathbb{Q}(\zeta_{3},\sqrt{6})$ and one computes that $\Delta_{K^{+}}=24$ and $\Delta_{K}=24^{2}$. In both cases $K/K^{+}$ is unramified.
Hence, $f(t)$ must be ordinary. Then, by Lemma 1.12, either $b=1-2q$ or, $b=2-2q$ and $q$ is odd. In the first case, Equation (3) implies that $N=1$, which cannot happen by assumption. So we must be in the second case, that is, iv holds. Equation (3) implies that if $b=2-2q$ (with $q$ odd) then $N\in\{1,2,4\}$, and in fact $N\in\{2,4\}$ since we are assuming that $K/K^{+}$ is ramified. Since $\mathfrak{m}$ is the unique maximal ideal of $\mathcal{O}^{+}$ above $2$ by Lemma 3.1, we obtain that $\mathfrak{m}$ ramifies in $K$, that is, ii holds.
[…]
@@
If $K/K^{+}$ is ramified or if there exists a prime in $K^{+}$ that divides $(\pi-q/\pi)$ and which stays inert in $K$, then, by [14, Thm. 1.1], there is a principally polarised abelian surface in the isogeny class associated to $f(t)$. Hence $f\in\mathcal{P}^{\mathrm{irr}}_{\mathrm{Wres}}$ by Theorem 1.4.
@@
If $f(t)$ is ordinary then [13, proof of Lemma 12.2] shows that $K$ contains $M=\mathbb{Q}(\zeta_{3})$. If $f(t)$ is not ordinary, then [15, Thm. 2] tells us that
@@
Then $\zeta_{3}\mapsto-\pi^{2}/q$ gives an embedding of $M$ in $K$ (cf. [15, Sec. 4]). ∎
@@
Let $f(t)\in\mathcal{P}^{\mathrm{irr}}_{\mathrm{npp}}\sqcup\mathcal{P}^{\mathrm{irr}}_{\mathrm{Wres}}$. The possible factorisations of $2$ in $K$ are as follows and all of them do occur.
@@
By Lemma 1.14, $2$ is not inert in $K$. We now show that $2$ cannot be totally split in $K$. If $f(t)\in\mathcal{P}^{\mathrm{irr}}_{\mathrm{Wres}}$ then $2$ ramifies in $K^{+}$ by Lemma 3.1 and we are done. Assume then that $f(t)\in\mathcal{P}^{\mathrm{irr}}_{\mathrm{npp}}$. Then, by Lemma 3.4, $K$ contains $M=\mathbb{Q}(\zeta_{3})$. Note that $2$ is inert in $M$, so $2$ is not totally split in $K$.
@@
We start by proving i. So, assume that $f(t)\in\mathcal{P}^{\mathrm{irr}}_{\mathrm{npp}}$. Then $M=\mathbb{Q}(\zeta_{3})$ is a subfield of $K$ by Lemma 3.4. Since $2$ is inert in $M$ then $2$ cannot be totally ramified in $K$. Also, $K/K^{+}$ is unramified by Lemma 3.3. If $2$ is inert in $K^{+}$ then we must have $2\mathcal{O}=\mathfrak{M}_{1}\mathfrak{M}_{2}$, with $\mathfrak{M}_{1}=\bar{\mathfrak{M}}_{2}$. If $2$ splits in $K^{+}$ then $2\mathcal{O}=\mathfrak{m}_{1}\mathfrak{m}_{2}$, that is each maximal ideal of $O^{+}$ above $2$ stays inert in $K$. In particular, $\mathfrak{M}_{1}=\overline{\mathfrak{M}}_{1}$ and $\mathfrak{M}_{2}=\overline{\mathfrak{M}}_{2}$. If $2$ is ramified in $K^{+}$ then $2\mathcal{O}=\mathfrak{M}^{2}$ must occur and $\mathfrak{M}$ is stable under conjugation. Part ii is Lemma 3.1. A search on the Weil polynomials shows that all listed factorisation occur; see Example 6.4. ∎
@@ 4. Kernels of polarisations of abelian varieties
$\mathcal{B}(S)=\frac{Z(S)}{(B(S)\cdot\mathop{\rm Pr}\nolimits_{S}(K^{\dagger}))},$
@@
    $f^{+}(t)\bmod\ell$ has a linear factor.
@@
We now show that i implies iii. So, assume also that $M$ as above satisfies $|M|=\ell^{2}$ for a rational prime $\ell$. Then $r=1$, $n_{1}=1$ and $R^{+}/\mathfrak{n}_{1}\simeq\mathbb{F}_{\ell}$. This implies that $f^{+}(t)\bmod\ell$ has a linear factor (cf. [26, Thm. 8.2]).
@@
This is part of [14, Prop. 6.2]. ∎
@@
This is a special case of [14, Prop. 6.4]. ∎
@@
This is part of [14, Prop. 7.1]. ∎
@@ 5. Polarisations of degree $4$ on abelian surfaces
In Section 2, we showed that an abelian surface with Weil polynomial $f\in\mathcal{P}^{\mathrm{irr}}_{\mathrm{npp}}\sqcup\mathcal{P}^{\mathrm{irr}}_{\mathrm{Wres}}$ contains an $\mathbb{F}_{q}$-irreducible curve of arithmetic genus $3$ if and only if it admits a polarisation with kernel of order $4$. We now apply the results developed in Section 4 to characterise isogeny classes defined by a polynomial in $\mathcal{P}^{\mathrm{irr}}_{\mathrm{npp}}\sqcup\mathcal{P}^{\mathrm{irr}}_{\mathrm{Wres}}$ not containing abelian surfaces with an $\mathbb{F}_{q}$-irreducible curve of genus $3$ lying on.
@@
Since $f(t)\in\mathcal{P}^{\mathrm{irr}}_{\mathrm{Wres}}$, the obstruction element $I_{\mathcal{C}}$ is trivial by Theorem 4.10. Hence, there is an $R$-module $M$ of order $4$ which is attainable if and only if the image $z_{\mathfrak{l}}$ of $[R/\mathfrak{l}R]_{R}$ in $\mathcal{B}(R)$ is trivial. Note that
@@
In Theorem 5.7, we wrap-up the discussion above. Moreover, in certain cases with $\mathcal{C}$ ordinary, we show that if there is an abelian variety with polarisation with kernel of order $4$ then there is an abelian variety with polarisation with kernel of order $4$ and maximal endomorphism ring. To do so, we use results from [13]. The construction of $\mathcal{B}(\mathcal{O})$ in [13] is slightly different form the one in [14] that we have presented above, but certainly equivalent in view of Proposition 4.8 and [13, Prop. 10.1].
@@
Clearly strongly attainable implies attainable, while the contrary does not hold in general. A characterisation of strongly attainable $R$-modules is given in [13, Prop. 5.7], repeated below as Proposition 5.5 for convenience. Such characterisation relates modules with a particular element $I_{K,\Phi}$ of $\mathcal{B}(\mathcal{O})$ which depends on a choice of a CM-type $\Phi$ of $K$. The definition of $I_{K,\Phi}$ is given in [13, Def. 5.2]. For our purposes, it sufficient to know whether it is trivial or not in $\mathcal{B}(\mathcal{O})$.
@@
Assume that $f\in\mathcal{P}^{\mathrm{irr}}_{\mathrm{npp}}$ is ordinary. Then $\mathcal{C}$ does not contain a principally polarised abelian variety. The result follows from [13, Prop. 11.3] and [13, Cor. 11.4].
Assume from now that $f\in\mathcal{P}^{\mathrm{irr}}_{\mathrm{Wres}}$ is ordinary. If $K/K^{+}$ is ramified then the $I_{K,\Phi}=0$ by [13, Prop. 11.1]. If $K/K^{+}$ is unramified and there is no maximal ideal of $\mathcal{O}^{+}$ dividing $\pi-q/\pi$ which stays inert in $K/K^{+}$ then the result follows, again, by combining [13, Prop. 11.3] and [13, Cor. 11.4] and the observation that $\mathcal{C}$ contains a principally polarised abelian variety. ∎
@@
[13, Prop. 5.7] Let $M$ be finite $R$-module. Then $M$ is strongly attainable if and only if $M\simeq\mathcal{O}/\mathfrak{a}\mathcal{O}$ for some $\mathcal{O}^{+}$-ideal $\mathfrak{a}\subseteq\mathcal{O}^{+}$ such that the image in $\mathcal{B}(\mathcal{O})$ of the class of $\mathcal{O}^{+}/\mathfrak{a}$ in $G(\mathop{\rm Mod}\nolimits_{\mathcal{O}^{+}})$ equals $I_{K,\Phi}$.
@@
The next theorem completes the proof of Main Theorem 3.
@@
Assume that $f\in\mathcal{P}^{\mathrm{irr}}_{\mathrm{npp}}\sqcup\mathcal{P}^{\mathrm{irr}}_{\mathrm{Wres}}$ and let $\mathcal{C}$ be the corresponding isogeny class.
@@
In the case when $f(t)\in\mathcal{P}^{\mathrm{irr}}_{\mathrm{Wres}}$, we completely characterise when the equivalent conditions ia,ib and iiif hold in terms of the coefficients of $f(t)$ in iiig. It is easy to obtain a characterisation in terms of the coefficients of $f(t)=t^{4}+at^{3}+bt^{2}+aqt+q^{2}$ also when $f(t)\in\mathcal{P}^{\mathrm{irr}}_{\mathrm{npp}}$. Indeed, if we write $\Delta_{f^{+}}=a^{2}-4(b-2q)=c^{2}d$ for integers $c$ and $d$ with $d$ squarefree, then it is well known that iie holds if and only if $d\equiv 5\bmod 8$. Moreover, if $q$ is even (or equivalently $a$ is odd) then $f^{+}(t)\equiv t^{2}+t+1\bmod 2$ is irreducible, which implies that $2$ is inert in $K^{+}$. Similarly, if $a$ is even and $a+b\not\equiv 1\bmod 4$ then $f^{+}(t)\equiv(t+1)^{2}$ and the remainder of the division of $f^{+}(t)$ by $t+1$ is not divisible by $4$. Hence, $2$ ramifies in $K^{+}$ by the Kummer–Dedekind Theorem [26, Thm. 8.2].
Remark 5.9.
The statement Theorem 5.7.ii can be deduced from the claim contained in [12, page 155, 2nd par. section 3], stated without a detailed proof.